A generic salt, AB2, has a molar mass of 317 g/mol and a solubility of 4.30 g/L at 25 °C. What is the Ksp of this salt at 25 °C?
Answer #1

mass solubility of salt = 4.30 g/L

molar solubility of salt = (mass solubility of salt) / (molar mass of salt)

molar solubility of salt = (4.30 g/L) / (317 g/mol)

molar solubility of salt = 0.0136 M

Ksp AB2 = [A][2B]2

Ksp AB2 = (s) * (2s)2

Ksp AB2 = (s) * (4s2)

Ksp AB2 = 4s3

where s = solubility

Ksp AB2 = 4 * (0.0136 M)3

Ksp AB2 = 9.98 x 10-6